[Brahms-l] Shift report 20050527 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri May 27 2005 - 23:48:06 EDT
Shifter leader was ejk
Shifters were BK, PS


1. We had 3.5 beam hours.
   LINAC problem, and restored.
   11:46pm : lost beam

2. 14992-14998 : data run
   14999, 15000 : pedestal run

3. BB big tubes are not stable. 
   We managed to turn on BB HV.
   We need to check!
   

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri May 27 23:48:20 2005

This archive was generated by hypermail 2.1.8 : Fri May 27 2005 - 23:48:32 EDT